OpenStudy (anonymous):

Which function has the greatest rate of change? A. x y 3 3 4 6 5 9 B. y = 4x – 7 C. 12x + 6y = 18 D. A linear function that goes through point 1, negative 1 and point 0, 4

OpenStudy (anonymous):

@iGreen

OpenStudy (invinceble):

A

OpenStudy (anonymous):

No direct answers. @invinceble

OpenStudy (anonymous):

xD

OpenStudy (igreen):

For A, take two points and put them in the slope formula. Let's take (3, 3) and (4, 6). \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{6-3}{4-3}\) Subtract: \(m = \dfrac{3}{1}\) Can you divide that? @RainbowBubbles01

OpenStudy (anonymous):

its 3

OpenStudy (anonymous):

Yup.

OpenStudy (igreen):

Yep, so the slope for A is 3.

OpenStudy (igreen):

Can you tell what the slope for B is? y = 4x – 7

OpenStudy (igreen):

*y = 4x - 7

OpenStudy (anonymous):

4

OpenStudy (igreen):

Yep, so the slope for B is 4.

OpenStudy (anonymous):

o

OpenStudy (anonymous):

ok i meant

OpenStudy (igreen):

For C, we have to change it to slope-intercept form: 12x + 6y = 18 Subtract 12x to both sides: 6y = -12x + 18 Divide 6 to both sides: y = -2x + 3 What's the slope of C?

OpenStudy (anonymous):

-2

OpenStudy (igreen):

Yep.

OpenStudy (igreen):

Well the equation for D will be: y = -5x + 4 So what's the slope?

OpenStudy (anonymous):

-5

OpenStudy (igreen):

Yep, so: A is 3 B is 4 C is -2 D is -5 Which one is the greatest?

OpenStudy (anonymous):

b

OpenStudy (igreen):

Yep, so B is your answer.

OpenStudy (anonymous):

can i ask 2 more?

OpenStudy (anonymous):

Which function has the greatest starting value? A. x y 1 −3 2 −8 3 −13 B. y = 3x – 8 C. 8x + 4y = 16 D. A linear function that goes through point 2, 0 and point 0, 2